Sylvia pushes a box with a force of 400N. If the box moves with an acceleration of 1.2 m/s2, whats is its mass


1
Expert's answer
2021-02-10T10:07:07-0500
"F=ma,""m=\dfrac{F}{a}=\dfrac{400\ N}{1.2\ \dfrac{m}{s^2}}=333.3\ kg."